OpenStudy (anonymous):

The midpoint of EF M(4,10) . One endpoint is E(2,6) find the coordnates of the other endpoint F a.(6,14) b.(6,17) c.(10,17) d.(14,8)

OpenStudy (anonymous):

it a yall slow

OpenStudy (anonymous):

@LieutenantGeneral

OpenStudy (anonymous):

@shinalcantara you need to type faster we don't need the steps we need the answer

OpenStudy (anonymous):

ole shorty getting blocked !!!

OpenStudy (shinalcantara):

Use midpoint formula \[x_m = \frac{ x_2+x_1 }{ 2 }\] \[y_m = \frac{ y_2+y_1 }{ 2 }\] ------------------- Since you are given the coordinates for the midpoint and the other end then, for x-coordinate: \[4 = \frac{ 2+x_1 }{ 2 }\] \[4(2) = 2 + x_1\] \[x_1 = 8 - 2\] \[x_1 = 6\] ------------------- \[10 = \frac{ 6+y_1 }{ 2 }\] \[10(2) = 6 + y_1\] \[20 - 6 = y_1\] \[y_1 = 14\]

OpenStudy (shinalcantara):

@jimmiedeel then you answer that question! -_-

OpenStudy (anonymous):

I did I gave him the answer so he would be quiet unlike you who had to write a whole pharagraph just to answer one stupid thing

OpenStudy (shinalcantara):

@jimmiedeel have you ever read the terms and conditions of os? (-_-) We don't just give answers right away @iGreen i need help lol

OpenStudy (anonymous):

yall stop arging but thanks for the answer @shinalcantara

OpenStudy (anonymous):

Well in my code of conduct it says otherwise and sahvalielbrown goes to my school so mind your own business thanks, Jimmiedeel

OpenStudy (igreen):

@jimmiedeel Nobody cares about 'your code-of-conduct'. Your code-of-conduct doesn't matter, because this site doesn't belong to you. You're not supposed to give answers, you're supposed to guide the asker to the answer.